# Scene-Text Grounding for Text-Based Video Question Answering

### v1

@@ I Introduction
Based on our ViTXT-GQA dataset, we source the baseline methods from 3 groups: TextVQA [16], TextVideoQA [42], and Weakly-grounded VideoQA [13, 24]. For methods from image-based TextVQA and TextVideoQA, we conduct a post-hoc analysis of the attention distribution corresponding to the predictions. For methods from Weakly-grounded VideoQA, we analyze their grounding outputs derived from the respective grounding modules. Our findings reveal that existing models struggle to predict visually grounded answers, despite their relatively decent QA behavior. Moreover, our T2S-QA improves over these methods for both grounding and QA. Yet, the gap between our results and human performance is still large. For example, humans can correctly ground 77% of the answers to the questions, whereas T2S-QA only achieves 28%. In addition, we find that there is a clear discrepancy between grounding and QA performance for both human (e.g., 77% vs. 64%) and T2S-QA (28% vs. 10%). On the one hand, the higher grounding and much lower QA accuracy of T2S-QA indicate that incorrect scene text recognition still seriously jeopardizes the final QA performance. On the other hand, the lower QA accuracy of human confirms the ineffectiveness of exact word matching for answer evaluation, and suggests the advantage and importance of our grounding evaluation to directly accept scene texts as visual answers. Finally, the large discrepancy between our method and human performance reveals that the poor TextVideoQA results of existing methods are largely attributed to the challenge posed by optical character recognition (OCR) in the video domain.
@@
  • We present a cascaded temporal-to-spatial scene-text grounding and question-answering model, and design a disentangled temporal and spatial contrastive learning strategy for weakly-supervised scene-text grounding.
@@
  • We thoroughly analyze the model behaviors, revealing the severe deficiency of existing scene-text recognition systems and the challenge of evaluating scene-text answers.
Figure 2: Overview of our T2S-QA. It mainly consists of three components: 1) the Feature Representation prepares the features of the question $\mathrm{Q}$, video frames $\mathrm{F}$, and OCR tokens $\mathrm{O}$; 2) the Contrastive Temporal-Spatial Grounding adopts a two-stage fine-grained grounding approach; 3) the Answer Decoder integrates the grounded frames $\mathrm{F}^{+}$, the grounded OCR tokens $\mathrm{O}^{+}$, and the question $\mathrm{Q}$ to achieve answer generation. In the Optimization process, we introduce a contrastive learning mechanism to help improve the question answering and answer grounding capabilities of our model.
@@ II-A Text-Based Video Question Answering
TextVideoQA aims to answer questions by jointly reasoning about scene texts and visual contents in videos. Unlike TextVQA [16, 43, 15], TextVideoQA focuses on developing the ability of machines to answer questions based on scene texts in dynamic environments which is more practical. Zhao et al. [42] first identify this challenge and build a benchmark M4-ViteVQA, whose videos factor nine scenarios (i.e., shopping, traveling, driving, vlog, sport, advertisement, movie, game, and talking), and develop a baseline model T5-ViteVQA. Tom et al. [31] propose the RoadTextVQA dataset on driving videos, focusing on questions that require understanding scene texts on road signs. Jahagirdar et al.[17] introduce NewsVideoQA to study TextVideoQA in news videos.
@@
A key challenge faced by the VQA [8, 27, 24] community is the issue of language bias [20, 14, 25]. Specifically, models can answer questions by exploiting language priors in datasets without considering the visual content. To some extent, this problem is difficult to detect as commonly used evaluation metrics solely take account into textual answers to assess model performance. In image-based VQA tasks, many works [4, 5, 29, 12] propose new benchmarks for answer grounding to support models to localize visual evidence that humans rely on to answer visual questions. This is due to the observation that answer grounding can serve as a valuable foundation for debugging VQA models. However, related research is relatively scarce in VideoQA. TVQA [21] and TVQA+ [22] pioneer in study grounded VideoQA, yet in full supervision. NExT-GQA [38] firstly studies weakly-supervised grounding for VideoQA, yet focusing on temporal dimension. Moreover, all of they concentrate on grounding the normal visual objects or video moments. In contrast, we focus on spatio-temporally grounding the scene text regions, with our techniques also centering on weakly-supervised grounding.
@@
In VideoQA task [37, 33, 6], several methods [28, 36, 38, 39, 13, 24] introduce weakly-supervised video grounding techniques to improve the QA performance and model interpretability [44]. These methods locate critical frames [28, 38, 39] and visual objects [13, 24] to aid question answering. Specifically, for video temporal grounding, [28] and [39] select the temporal proposal with the highest similarity to the language query. Besides, [38] designs a Gaussian mask learning module to ground a temporal segment.
Compared with VideoQA, TextVideoQA focuses on understanding local and even tiny scene texts in the video, where frame-level grounding struggles to capture subtle scene text details. In this work, we study spatio-temporal grounding [41] for the proposed Grounded TextVideoQA task. For video spatio-temporal grounding, MIST [13] designs an iterative attention module to select video segment and region features and fuse them for question answering. TranSTR [24] locates key video frames and objects and uses object-enhanced frame features for QA. However, these methods rely on the similarity between frame (object) features and questions to select critical elements, lacking effective self-supervision to constrain the selection during training. To address this, we design a disentangled temporal and spatial contrastive learning strategy that distinguishes positive and negative video frames and scene text regions in the video and enlarges the semantic difference between positive and negative features. This approach injects strong self-supervision signals into the model to distinguish question-related video frames and scene texts, thereby effectively achieving scene text grounding.
@@ III Method
Given a video and a question, Grounded TextVideoQA aims to encapsulate visual content and question semantics, and then generate textual answers based on the localized question-related scene texts from the video. Meanwhile, the localized scene text is used as an auxiliary visual answer to support the interpretability of the textual answer. As shown in Fig. 2, we design a cascaded temporal-to-spatial grounding architecture (T2S-QA) to perform answer grounding and grounded QA. Here, we provide a detailed introduction to each component.
@@ III-A Feature Representation
Video features. Given a video, we uniformly sample $T$ frames and extract frame feature $F=\{f_{t}\}^{T}_{t=1}\in\mathbb{R}^{T\times d}$ to represent a video. For the $i$-th video frame, its features include two parts: 1) visual feature $f^{v}_{i}\in\mathbb{R}^{1024}$ extracted by a pre-trained visual backbone which is frozen during training; 2) temporal id feature $f^{t}_{i}\in\mathbb{R}^{50}$ where the temporal id $i$ is assigned to denote the $i$-th video frame and encoded by a embedding layer as a temporal information. We merge two features to obtain the final video frame feature.
$f_{i}=\mathrm{LN}({W_{1}}f^{v}_{i}+{W_{2}}f^{t}_{i}),$ (1)
@@
OCR features. Following [42], we employ a scene text detection system [35] to obtain bounding boxes and tracking ids of the scene texts in each video frame, and then use a scene text recognition system [11] to read the scene texts according to their bounding boxes. (For simplicity, the scene text detection system and recognition system are collectively referred to as OCR systems.) After that, we obtain a set of OCR tokens $O=\{o_{t,s}\}^{T,S}_{t=1,s=1}\in\mathbb{R}^{T\times S\times d}$ in the video, where $T$ is the frame number and $S$ is the OCR token number in each frame. For the $j$-th OCR token in the $i$-th frame, its features consist of five aspects: 1) FastText feature $o^{f}_{i,j}\in\mathbb{R}^{300}$ [3] providing essential word information; 2) pyramidal histogram of characters (PHOC) feature $o^{p}_{i,j}\in\mathbb{R}^{604}$ [1] representing character features; 3) normalized bounding box feature $o^{b}_{i,j}\in\mathbb{R}^{4}$ as the visual geometric features; 4) temporal id feature $o^{t}_{i,j}\in\mathbb{R}^{50}$ to provides temporal features consistent with the video frames; 5) track id feature $o^{a}_{i,j}\in\mathbb{R}^{50}$ as the identification features. To project the representations into a common $d$-dimensional space, we apply a linear mapping on the above features respectively and sum them up to get the final representation of the OCR token features.
$o_{i,j}\!$ $=\!\mathrm{LN}({W_{3}}o^{f}_{i,j}\!+\!{W_{4}}o^{p}_{i,j}\!+\!{W_{5}}o^{t}_{i,j}\!+\!{W_{6}}o^{a}_{i,j})\!+\!\mathrm{LN}({W_{7}}o^{b}_{i,j}),$ (2)
@@ III-B1 Temporal Grounding
In most cases, scene text used to answer questions tends to appear in a portion of video frames according to Fig. 3 (b). To identify the critical frames from the video, TG takes the video frames $\mathrm{F}$ and question $\mathrm{Q}$ as input and adaptively selects question-relevant frames. It first takes a self-attention layer in Eqn. 4 to learn a global question embedding $\mathrm{q}_{g}\in\mathbb{R}^{d\times 1}$. Then, a cross-attention is applied to represent the correlation probability between each video frame and the question. In Eqn. (5) and (6), it respectively calculates two attention matrixes $\mathrm{f}^{+}\in\mathbb{R}^{T\times 1}$ and $\mathrm{f}^{-}\in\mathbb{R}^{T\times 1}$ separately to learn the distributions of positive frames and negative frames in the video. Following [23], we leverage Gumbel-Softmax [18] to manage a differentiable discrete selection between the two calculated attention probabilities and obtain a contrastive map $\mathrm{M}_{f}\in\mathbb{R}^{T\times 2}$ to distinguish between positive and negative frames in the video.
@@
The scene text in the video may be intermittently blocked or blurred during movement. Therefore, not all positive video frames related to the question can locate clear and unobstructed scene text to answer the question. We need to further filter the positive and negative video frames based on their similarity to the given question. Specifically, for positive frames, we utilize $\textrm{Top}_{K}^{\uparrow}(\cdot)$ function to select $K_{1}$ video frames with the highest attention weight from $\mathrm{f}^{+}$ and gather their corresponding frame features as $\mathrm{{F}^{+}}\in\mathbb{R}^{K_{1}\times d}$. For negative frames, we take $\textrm{Top}_{K}^{\downarrow}(\cdot)$ function to select $K_{1}$ video frames with the lowest attention weight from $\mathrm{f}^{-}$ and gather their corresponding frame features as $\mathrm{{F}^{-}}\in\mathbb{R}^{K_{1}\times d}$.
@@ III-B2 Spatial Grounding
Our analysis reveal that answers only occupy a very small area in the video frame, as shown in Fig. 3 (c). Therefore, we first need to distinguish the scene texts related to the question from numerous elements, and then locate a few critical strongly associated scene texts as visual answers. To achieve that, SG facilitates adaptive OCR token selection similar to Eqn. (4) $\sim$ (9). Concretely, for the $i$-th critical frame, SG is fed with $S$ object features detected on the frame, enabling it to differentiate positive OCR tokens from negative ones based on their similarity to the given question. For positive OCR tokens, we select $K_{2}$ OCR tokens in each frame with the highest attention weight from $\mathrm{o}^{+}$ and gather their corresponding OCR features as $\mathrm{{O}^{+}}\in\mathbb{R}^{K_{1}\times K_{2}\times d}$. For negative OCR tokens, we select $K_{2}$ OCR tokens in each frame with the lowest attention weight from $\mathrm{o}^{-}$ and gather their corresponding OCR features as $\mathrm{{O}^{-}}\in\mathbb{R}^{K_{1}\times K_{2}\times d}$. It is worth noting that the detected OCR results vary for each frame, and SG is applied independently to each frame. Hence, even if we use hyperparameter $K_{2}$ to select OCR tokens in each frame, the numbers of critical OCR tokens in each frame are different.
Up to now, we have obtained the pair of the $K_{1}$ positive video frames and $K_{1}\times K_{2}$ positive OCR tokens, namely $\{\mathrm{{F}^{+}},\mathrm{{O}^{+}}\}$, and the pair of the $K_{1}$ negative video frames and $K_{1}\times K_{2}$ negative OCR tokens, namely $\{\mathrm{{F}^{-}},\mathrm{{O}^{-}}\}$. Notably, we take the bounding boxes of $\mathrm{{O}^{+}}$ as grounded OCR tokens for answer grounding evaluation, the evaluation details are elaborated in Sec. V.

### v2

@@ I Introduction
Based on our ViTXT-GQA dataset, we source the baseline methods from three groups: TextVQA [20], TextVideoQA [55], and weakly-grounded VideoQA [14, 29]. For methods from image-based TextVQA and TextVideoQA, we conduct a post-hoc analysis of the attention distribution corresponding to the predictions. For methods designed for weakly-grounded VideoQA, we analyze their grounding outputs derived from the respective grounding modules. Our findings reveal that existing models struggle to predict visually grounded answers, despite their relatively decent QA behavior. The results show our T2S-QA improves over these methods for both grounding and QA. Additionally, we evaluate two advanced multimodal large language models (MLLMs) on ViTXT-GQA: the open-source Qwen2-VL [44] and the closed-source GPT-4o-mini [34]. The experimental results show that T2S-QA has better scene text grounding capabilities than them, indicating the huge room for improvement of current MLLMs in scene-text grounding. Moreover, we find the performance gap between our T2S-QA and human is still large. For example, humans can correctly ground 77% of the answers to the questions, whereas T2S-QA only achieves 28%. In addition, we find that there is a clear discrepancy between grounding and QA performance for both human (77% vs. 64%) and T2S-QA (28% vs. 10%). On the one hand, the higher grounding and much lower QA accuracy of T2S-QA indicate that incorrect scene text recognition still seriously jeopardizes the final QA performance. On the other hand, the lower QA accuracy of human confirms the ineffectiveness of exact word matching for answer evaluation, and suggests the advantage and importance of our grounding evaluation to directly accept scene texts as visual answers. Finally, the large discrepancy between T2S-QA and human reveals that the poor TextVideoQA results of existing methods are largely attributed to the challenge posed by scene-text detection and recognition in the video domain.
@@
  • We propose a cascaded temporal-to-spatial scene-text grounding and question-answering model, T2S-QA, with a disentangled temporal-spatial contrastive learning strategy for weakly-supervised scene-text grounding.
@@
  • We thoroughly analyze the model behaviors, revealing significant challenges in performing Grounded TextVideoQA.
Figure 2: Overview of our T2S-QA model. It mainly consists of three components: (1) the Feature Representation prepares the features of the question $\mathrm{Q}$, video frames $\mathrm{F}$, and OCR tokens $\mathrm{O}$; (2) the Contrastive Temporal-Spatial Grounding adopts a two-stage fine-grained grounding approach; (3) the Answer Decoder integrates the grounded frames $\mathrm{F}^{+}$, the grounded OCR tokens $\mathrm{O}^{+}$, and the question $\mathrm{Q}$ to achieve answer generation. In the Optimization process, we introduce a contrastive learning mechanism to help improve the question answering and answer grounding capabilities of our model.
@@ II-A Text-Based Video Question Answering
TextVideoQA aims to answer questions by jointly reasoning about scene texts and visual contents in videos. Unlike TextVQA [20, 56, 18, 57], TextVideoQA [55, 41, 21, 58] focuses on developing the ability of machines to answer questions based on scene texts in dynamic environments. Zhao et al. [55] first identify this challenge and build a benchmark M4-ViteVQA, whose videos factor nine scenarios (i.e., shopping, traveling, driving, vlog, sport, advertisement, movie, game, and talking), and develop a baseline model T5-ViteVQA. Tom et al. [41] propose the RoadTextVQA dataset on driving videos, focusing on questions that require understanding scene texts on road signs. Jahagirdar et al.[21] introduce NewsVideoQA to study TextVideoQA in news videos.
@@
A key challenge faced by the VQA [38, 9, 35, 29] community is the issue of language bias [23, 40, 17, 32]. Specifically, models can answer questions by exploiting language priors in datasets without considering the visual content. To some extent, this problem is difficult to detect as commonly used evaluation metrics solely take account into textual answers to assess model performance. In image-based VQA tasks, many works [4, 5, 37, 13] propose new benchmarks for answer grounding to support models to localize visual evidence that humans rely on to answer visual questions. This is due to the observation that answer grounding can serve as a valuable foundation for debugging VQA models. However, related research is relatively scarce in VideoQA. TVQA [24] and TVQA+ [25] pioneer in study grounded VideoQA, yet in full supervision. NExT-GQA [49] firstly studies weakly-supervised grounding for VideoQA, yet focusing on temporal dimension. Moreover, all of them concentrate on grounding the normal visual objects or video moments. In contrast, we focus on spatio-temporally grounding the scene text regions, with our techniques also centering on weakly-supervised grounding.
@@
In VideoQA task [51, 48, 43, 7], several methods [36, 47, 49, 52, 14, 29] introduce weakly-supervised video grounding techniques to improve the QA performance and model interpretability [59]. These methods locate critical frames [36, 27, 28, 52] and visual objects [14, 29, 50] to aid question answering. For example, [36] and [52] select the temporal proposal with the highest similarity to the language query. [49] design a Gaussian mask learning module to ground temporal segments for answer prediction.
Compared with VideoQA, TextVideoQA focuses on understanding local and even tiny scene texts in the video, where frame-level grounding struggles to capture subtle scene text details. In this work, we study spatio-temporal grounding [54] for the proposed Grounded TextVideoQA task. For video spatio-temporal grounding, MIST [14] designs an iterative attention module to select video segment and region features and fuse them for question answering. TranSTR [29] locates key video frames and objects and uses object-enhanced frame features for QA. However, these methods rely on the similarity between frame (object) features and questions to select critical elements, lacking effective self-supervision to constrain the selection during training. To address this, inspired by disentangled learning approaches [31] and contrastive learning strategy [30, 16, 15], we design a disentangled temporal and spatial contrastive learning strategy that distinguishes positive and negative video frames and scene text regions in the video and enlarges the semantic difference between positive and negative features. This approach injects strong self-supervision signals into the model to distinguish question-related video frames and scene texts, thereby effectively achieving scene text grounding. Our method distinguishes positive and negative video frames and scene text regions in the video and enlarges the semantic difference between positive and negative features. This approach injects strong self-supervision signals into the model to distinguish question-related video frames and scene texts, thereby effectively achieving scene text grounding.
@@ III Method
Given a video and a question, Grounded TextVideoQA aims to encapsulate visual content, scene text, question semantics, and generate textual answers. Meanwhile, the scene text should be explicitly localized to substantiate the predicted textual answer or directly act as visual answer. To solve the task, localizing the question-related scene text is of crucial importance. As such, we propose a cascaded Temporal-to-Spatial Grounding and QA (T2S-QA) architecture as depicted in Fig. 2. T2S-QA  comprises three core modules: 1) Feature Representation module extracts features for the question, video frames, and scene texts; 2) the Contrastive Temporal-Spatial Grounding (TSG) module executes temporal grounding to identify question-relevant video frames, followed by spatial grounding to pinpoint critical scene text regions within the identified frames; 3) the Answer Decoder module integrates the grounded video frames, scene text, and question to decode the final answer. For optimization, we curate positive and negative pairs to optimize the alignment between scene texts and the question spatio-temporally.
@@ III-A Feature Representation
Video features. Given a video, we uniformly sample $T$ frames $V=\{v_{t}\}^{T}_{v=1}$ and extract frame feature $F=\{f_{t}\}^{T}_{t=1}\in\mathbb{R}^{T\times d}$ to represent a video. For the $t$-th video frame, its features include two parts: 1) visual feature $f^{v}_{t}\in\mathbb{R}^{1024}$ extracted by a pre-trained visual backbone which is frozen during training; 2) temporal id feature $f^{te}_{t}\in\mathbb{R}^{50}$ where the temporal id $t$ is assigned to denote the $t$-th video frame and encoded by a embedding layer as a temporal information. We merge two features to obtain the final video frame feature.
$f_{i}=\mathrm{LN}({W_{1}}f^{v}_{t}+{W_{2}}f^{te}_{t}),$ (1)
@@
OCR features. Given the sampled video frames $V=\{v_{t}\}^{T}_{t=1}$, following [55], we employ a scene text detection system [46] to obtain bounding boxes and track ids of the scene texts in each video frame, and then use a scene text recognition system [12] to read the scene texts according to their bounding boxes. For simplicity, the scene text detection system and recognition system are collectively referred to as optical character recognition (OCR) systems. Finally, the OCR token set for a video is denoted as:
$\{\text{B}_{t},\text{Ta}_{t}\}^{T}_{t=1}=\text{Detector}(v_{t}),$ (2)
$\{r_{t,s}\}^{S}_{s=1}=\text{Recognizer}(v_{t},B_{t}),$ (3)
$o_{t,s}=\{r_{t,s},o^{b}_{t,s},o^{ta}_{t,s}\}^{T,S}_{t=1,s=1},$ (4)
where $\text{Detector}(\cdot)$ is the scene text detection system applied to frame $v_{t}$. $\text{Recognizer}(\cdot)$ is the scene text recognition system applied to frame $v_{t}$ and its bounding boxes ${\text{B}_{t}}$, where $B_{t}=\{o^{b}_{t,1},o^{b}_{t,2},\dots,o^{b}_{t,S}\}$ is the set of bounding boxes in frame $t$ and $S$ is the OCR token number in each frame, where each box contains the coordinates $[x_{t,s}^{tl},y_{t,s}^{tl},x_{t,s}^{br},y_{t,s}^{br}]$ for the top-left and bottom-right corners. $\text{Ta}_{t}=\{o^{ta}_{t,1},o^{ta}_{t,2},\dots,o^{ta}_{t,S}\}$ is set of track ids for the detected text regions in frame $t$. $r_{t,s}$ is the recognized scene text content.
After that, we obtain OCR features $O=\{o_{t,s}\}^{T,S}_{t=1,s=1}\in\mathbb{R}^{T\times S\times d}$ in the video. For the $s$-th OCR token in the $t$-th frame, its features consist of five aspects: 1) FastText feature $o^{f}_{t,s}\in\mathbb{R}^{300}$ [3] providing essential word information based on the recognized scene text content $r_{t,s}$; 2) pyramidal histogram of characters (PHOC) feature $o^{p}_{t,s}\in\mathbb{R}^{604}$ [1] representing character features based on the recognized scene text content $r_{t,s}$; 3) normalized bounding box feature $o^{b}_{t,s}\in\mathbb{R}^{4}$ as the visual geometric features; 4) temporal id feature $o^{te}_{t,s}\in\mathbb{R}^{50}$ to provides temporal features consistent with the video frames; 5) track id feature $o^{ta}_{t,s}\in\mathbb{R}^{50}$ as the identification features. To project the representations into a common $d$-dimensional space, we apply a linear mapping on the above features respectively and sum them up to get the final representation of the OCR token features.
$o_{t,s}\!$ $=\!\mathrm{LN}({W_{3}}o^{f}_{t,s}\!+\!{W_{4}}o^{p}_{t,s}\!+\!{W_{5}}o^{te}_{t,s}\!+\!{W_{6}}o^{ta}_{t,s})\!+\!\mathrm{LN}({W_{7}}o^{b}_{t,s}),$ (5)
@@ III-B1 Temporal Grounding
In most cases, scene text relevant to questions often appears in specific video frames as shown in Fig. 3. To align video content with question semantics and ground key video frames as context for effective question answering, we design a multi-step approach as follows.
Global Question Embedding. A global sentence representation, $\mathrm{q}_{g}$, is derived from word embeddings $\mathrm{q}_{l}$ using self-attention, capturing overall question semantics by emphasizing keywords.
@@
Correlation Scoring. The correlation-based softmax operation computes the relevance distribution between the global question embedding and the video frame features. Guided by the contrastive learning objective in Eqn. (14), we calculate the cross-attention scores for the positive frame distribution $\mathrm{f}^{+}\in\mathbb{R}^{T\times 1}$ and the negative frame distribution $\mathrm{f}^{-}\in\mathbb{R}^{T\times 1}$, respectively. This approach explicitly distinguishes between positive (relevant) and negative (irrelevant) frames, ensuring that the positive frames align with the question semantics. The contrastive learning objective further amplifies the semantic differences between positive and negative frames, enhancing the capture of question-relevant video information.
@@
Contrastive Mapping. To identify video frames that are relevant or irrelevant to the question, we apply the Gumbel-Softmax function to obtain the distribution of positive ($\mathrm{f}^{+}$) and negative ($\mathrm{f}^{-}$) frames. This operation enables a discrete distinction, ensuring that the positive frames align with the question semantics.
@@
Frame Filtering. Not all positive video frames carry clear and unobstructed scene text to answer the question. We need to further filter the positive and negative video frames based on their relevance to the given question. Specifically, we utilize Eqns (11) and (12) to enhance the semantic distinction between positive and negative frame features. Leveraging the cross-correlation scores computed by Eqns (8) and (9), we identify the top $K_{1}$ video frames from the positive (negative) frame set that exhibit the highest (lowest) relevance to the question. The obtained frame sets $\{\mathrm{F}^{+},\mathrm{F}^{-}\}$ demonstrate significant semantic divergence, quantified by their relevance to the question. Guided by the contrastive learning objective, the model is optimized to further amplify this divergence, enabling grounding of question-relevant video frames $\mathrm{F}^{+}$.
@@
Up to now, we have introduced how to ground the frames $\mathrm{F}^{+}$ that carry the key scene text for answering questions. These frames will be inputted into the answer decoder (Sec. III-C1) to serve as context for answer prediction.
@@ III-B2 Spatial Grounding
Our preliminary analysis reveals that most answers only occupy a very small area in the video frame, as shown in Fig. 3. Therefore, we first need to distinguish the scene texts related to the question from numerous irrelevant visual elements, and then locate a few critical strongly associated scene texts as visual answers. To achieve that, SG facilitates adaptive OCR token selection similar to Eqn. (7) $\sim$ (12). Concretely, for the $i$-th critical frame, SG is fed with $S$ object features detected on the frame, enabling it to differentiate positive OCR tokens from negative ones based on their similarity to the given question. For positive OCR tokens, we select $K_{2}$ OCR tokens in each frame with the highest attention weight from $\mathrm{o}^{+}$ and gather their corresponding OCR features as $\mathrm{{O}^{+}}\in\mathbb{R}^{K_{1}\times K_{2}\times d}$. For negative OCR tokens, we select $K_{2}$ OCR tokens in each frame with the lowest attention weight from $\mathrm{o}^{-}$ and gather their corresponding OCR features as $\mathrm{{O}^{-}}\in\mathbb{R}^{K_{1}\times K_{2}\times d}$. It is worth noting that the detected OCR results vary for each frame, and SG is applied independently to each frame. Hence, even if we use hyperparameter $K_{2}$ to select OCR tokens in each frame, the numbers of critical OCR tokens in each frame are different.
Up to now, we have obtained the pair of the $K_{1}$ positive video frames and $K_{1}\times K_{2}$ positive OCR tokens, namely $\{\mathrm{{F}^{+}},\mathrm{{O}^{+}}\}$, and the pair of the $K_{1}$ negative video frames and $K_{1}\times K_{2}$ negative OCR tokens, namely $\{\mathrm{{F}^{-}},\mathrm{{O}^{-}}\}$. Notably, we take the bounding boxes of $\mathrm{{O}^{+}}$ as grounded OCR tokens for answer grounding evaluation, the evaluation details are elaborated in Sec. V-A.